OpenStudy (anonymous):

How do i solve this equation? x^3+x^2-3x-3=0 Also how do i find the number and type of roots?

OpenStudy (anonymous):

please help me!!!!

OpenStudy (sosa4954):

Factor the equation. (x^2)(x+1) - 3(x+1) = 0 Combine like terms. (x+1)((x^2)-3) = 0 So x+1 = 0 or (x^2) - 3 = 0. x = -1 or (x^2) = 3 x = -1 or ±√3